OpenStudy (anonymous):

How do I put this in my graphing calculator and get the answer: (-3.449, -11.899) f(x)= (x^2 - 5x)/(x+1), x<-1

OpenStudy (anonymous):

\[f(x)=\frac{ x^2-5x }{ x+1 }, x<-1\]

hero (hero):

More than one point is the solution to that, so what is the original question associated with that function? Otherwise, no one will have a clue how to get that answer. Help us help you

OpenStudy (anonymous):

Find the maximum of this function

hero (hero):

To find the max value of the function 1. Find f'(x). 2. Set f'(x) = 0. 3. Solve for x using your calc. 4. Plug x back into the expression for f(x) 5. Evaluate the expression using your calc.

OpenStudy (anonymous):

for the derivative i got \[\frac{ x^2 +2x - 5 }{ (x+1)^2 } = 0\]

OpenStudy (anonymous):

so in order to find the zeros i would have to get the quadratic formula for this equation?

hero (hero):

Make sure to multiply both sides of the equation by (x+1)^2 first x^2 + 2x - 5 = 0

OpenStudy (anonymous):

so when i do that i'll just be left off with x^2 +2x - 5, then i use the quadratic formula to find the zeros?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

how did u get the x^2 +2x +4 part?

hero (hero):

Actually, hang on. I made a mistake

OpenStudy (anonymous):

"Find the maximum to these functions" the answer my teacher gave me was (-3.449, -11.899)

OpenStudy (anonymous):

he used the graphing calculator to get the answer but when I used the graphing calc. I got a way different answer than he did x-x

hero (hero):

Bro, I just told you I made a mistake. Will you let me correct it? PLEASE

hero (hero):

It was not necessary to repeat THE SAME THING TWICE

OpenStudy (anonymous):

sorry

hero (hero):

x^2 + 2x = 5 x^2 + 2x + 1 = 5 + 1 (x + 1)^2 = 6 x + 1 = ± √(6) x = ± √(6) - 1 x = √(6) - 1 x = -√(6) - 1

hero (hero):

There. That should work Obviously you will only be choosing the positive point

OpenStudy (anonymous):

how did uget the x^2 +2x +1 part

hero (hero):

You add (b/2)^2 to both sides. That's what is called completing the square.

hero (hero):

so since b = 2, (2/2)^2 = 1

hero (hero):

so you add 1 to both sides

hero (hero):

Sorry, you choose the negative x value, not the positive one

OpenStudy (anonymous):

so i dint have to use the quadratic formula for this?

hero (hero):

There's more than one way to solve quadratic equations. I suppose your teacher never told you that.

OpenStudy (anonymous):

well when i did the quadratic formula for this I got \[\frac{ -2 \pm \sqrt{24}}{ 2}\] is that correct?

hero (hero):

Yes, that is correct because if you keep reducing you end up with the same thing: \[\frac{-2 \pm \sqrt{(4)(6)}}{2}\] \[\frac{-2 \pm \sqrt{4}\sqrt{6}}{2}\] \[\frac{-2 \pm 2\sqrt{6}}{2}\] \[2\left(\frac{-1 \pm \sqrt{6}}{2}\right)\] \[-1 \pm \sqrt{6}\] \[\pm \sqrt{6} -1\]

OpenStudy (anonymous):

so i would always choose the answer that is negative for these kinds of problems? "Finding the maximum to these functions"

hero (hero):

Not always

hero (hero):

For this particular problem, it was stated that x < -1

OpenStudy (anonymous):

oh okay ty :DDD

OpenStudy (anonymous):

but my teacher used the graphing calculator for this, can u show me how he did b/c when I try to plug in the equation on my calc. i get a weird graph and can't find a maximum on there

hero (hero):

No, I cannot show you how he did it. I'm not your teacher and I don't have his graphing calc. All I can do is tell you how I would do it. Besides I use TI-Nspire Graphing Calc. Your teacher probably uses TI-89 or something older.

OpenStudy (anonymous):

oh. ok ty ^^

OpenStudy (anonymous):

can u hep me with another problem

hero (hero):

All you have to do is just post your question. Someone will help. But close this question first. Then post your question as a new one.